# Monte Carlo evaluations of gamma-ray and radio pulsar populations

### v1

@@ 4 Results
We evaluate different models based on two main approaches. First, we simply look at the number of Fermi-LAT detectable $\gamma-$ray pulsars predicted by each model and compare it with the known number of pulsars plus the number of PLUIDs in the 4FGL-DR4 Fermi-LAT catalog. Second, we look at the distribution of the modeled sample’s period ($P$), period derivative ($\dot{P}$), Galactic longitude, $(l)$ and latitude $(b)$ distributions to see how they compare with the observed sample. We do this as a three step qualitative process by first looking at the distribution in the $P-\dot{P}$ plot, then comparing the parameters’ cumulative distribution functions (CDFs) and finally evaluating their quantile-quantile (QQ) plots (Wilk & Gnanadesikan, 1968). For a given CDF, every probability value $p$ from the ordinate can be mapped to a quantile value $q(p)$ on the abscissa. For a particular parameter, plotting the $q(p)$ function for two different distributions (in this case, observed CDFs versus model CDFs of the parameter) on the same two-dimensional plot results in a QQ plot for that parameter. If both the distributions are identical, the resulting plot would be a straight line with a gradient of one, passing through the origin. More importantly, if either of the two distributions is a linear function of the other, the resulting QQ plot would still be a straight line albeit with a different slope and $y$-intercept. Any differences between the two distributions, specially in their tails, can be easily spotted on a QQ plot as deviations from linearity. These deviations are easier to spot then noticing subtle changes in the CDFs themselves which is why QQ plots were employed as a visual statistical tool in this work. The purpose of these investigations was not to discern if a model is better than another, but rather to check how the the underlying distribution resulting from each model differs from the observed one for a given parameter. We also show the $P-\dot{P}$ plot for the resulting populations from our radio synthesis compared with the observed sample which was used to run the Monte Carlo simulation in Fig. 3.
@@ 4.1 Millisecond pulsars (MSPs)
We show the number of Fermi-LAT detectable $\gamma-$ray pulsars predicted by each luminosity model in Fig. 2, along with the acceptable region described in Section 2 highlighted in green. We compare the observed and synthesized distribution for four of these parameters by looking at CDFs shown in Figs. 4, 5, 6 and 7. The observed sample contains 141 $\gamma-$ray MSPs which are in the Galactic plane (i.e., excluding globular cluster MSPs J1823--3021A and J1824--2452A) except for the $\dot{P}$ CDFs in 5. The observed sample in $\dot{P}$ CDFs contains 134 MSPs because 7 MSPs do not have a recorded $\dot{P}$ values in 3PC. Finally, we also provide a $P-\dot{P}$ plot for each of the models in Fig. 8. In this figure, each $P-\dot{P}$ plot has the same number of pulsars. This was done by randomly sampling 134 pulsars to match the number of known pulsars which have measured $\dot{P}$ and are shown in the same figure.
@@
Figure 4: CDFs showing the distribution of period for the observed and model populations.
Figure 5: CDFs showing the distribution of period derivative for the observed and model populations.
Figure 6: CDFs showing the distribution of Galactic longitude for the observed and model populations.
Figure 7: CDFs showing the distribution of Galactic latitude for the observed and model populations.
@@ 4.2 Canonical pulsars (CPs)
The 3D-plane suggests 440 Fermi-LAT detectable $\gamma-$ray sources, the 2D-plane suggest 257 of them and the heuristic model, which included beaming, suggests 157 of them. We show the $P-\dot{P}$ distribution for CPs compared with the observed sample in Fig. 9. Since the number of predicted CPs is more than the observed sample for all three models, we chose to show ony 150 randomly sampled CPs for each model in these plots. We also show the number of radio-loud $\gamma-$ray sources in this plot. We define radio-loud as the CP having a radio flux of greater than 100 $\mu$Jy.
@@ 5 Discussion
It is important to assess that the radio synthesis results in a population compatible with the observed radio pulsars. This can be done by qualitatively using the $P-\dot{P}$ plot shown in Fig. 3. We can see that the CP synthesis does well, as expected from the work of Johnston & Karastergiou (2017). Similarly, the MSP synthesized sample is able to cluster around the same point as the observed sample albeit with less dispersion. The main discrepancy between the model and observed pulsars in Fig. 3 is the lack of synthesized pulsars in the 20–100 ms period range. This is not surprising given that our modeling does not account for high-mass binary evolution which is thought to produce most of the pulsars in this region (i.e., the double neutron star binaries and the disrupted recycled pulsars; Tauris & van den Heuvel, 2023). These systems are not expected to have a significant impact on our main findings.
@@ 5.1 Millisecond pulsars (MSPs)
As can be seen in Fig. 2, only the log-normal, 2D-plane and the 3D-plane $\gamma-$ray luminosity model is in the acceptable region for number of Fermi-LAT detectable MSPs in the Galaxy. None of the other $\gamma-$ray luminosity models proposed in the literature so far are compatible with a population synthesis calibrated to the radio population. However, the observed sample size of PLUIDs we used was based on simple assumptions. More complicated methods which could arguably better predict the number of pulsars among this sample also exist in literature. These methods use classification from machine learning (ML) to untangle the nature of the unidentified sources. Zhu et al. (2024) use multiple ML models and then create a voting ensemble from those to classify unidentified sources (UIDs) after dividing the sample into a high Galactic latitude (HGL) region where $|b|>10^{\circ}$ and a low Galactic latitude (LGL) region where $|b|<10^{\circ}$. Their ensemble finds 88 PLUIDs in the HGL and 134 in the LGL regions respectively. The fact that they are unable to distinguish between MSPs and CPs makes a direct comparison with our work difficult. However, the vast majority of the HGL region pulsars are likely to be MSPs and the LGL region should have a CP majority. This suggests a lower limit of $88+134=222$ MSPs to be compatible with their work. All of our models satisfy this limit.
Looking at the $P-\dot{P}$ distributions in Fig. 8, we see that the observed sample clusters around the point at 3 ms and $10^{-20}$, respectively. This is not the case for the BPL and log-normal model as those pulsars have a lower $\dot{P}$s. The spread of the FP and GHF+18 models seem to be consistent with the observed sample. However, the FP models have a tighter spread than the GHF+18 models which better matches the observed sample. The observed sample does not contain any MSPs with $\dot{P}<10^{-21}$. The FP and GHF+18 models also avoid producing too many MSPs below this imaginary horizontal line on the plot.
The CDFs for these parameters help us understand the spread in these $P-\dot{P}$ plots in another way. In Fig. 4, we see that all models except BPL and Log-normal match the observed sample until period of 3 ms. 3D FP and OG model lack the required density in the 3–5 ms period interval. TPC and 2D FP are able to match the observed sample in the range but underdense in the 5–10 ms range. PSPC matches the observed distribution the closest. Similarly, the $P-\dot{P}$ distribution in Fig. 5 shows us that the BPL and log-normal models perform the worst. Overdensity in the $\dot{P}<10^{-20}$ s s-1 is seen for all models. 3D FP is able to best match the higher $\dot{P}$ range. The sky distributions for these simulated populations is shown in Figs. 6 ($l$) and 7 ($b$). The $l$ distributions are all in a reasonable agreement with the observed sample. Their noticeable features are that all models have a larger population under $l<0$ than the observed. The trend continues for $l>0$ for all models except the for 2D and 3D FP. The $b$ CDFs for all models except the GHF+18 models is also in close agreement with the observed distribution. The GHF+18 models are closer to the galactic plane than the observed sample.
We can find further insights into these distributions by looking at how each quantile of the parameter distributions compares with each other in a QQ plot. For MSPs, this is shown in Fig. 10. All $\gamma-$ray luminosity models except for LN are able to match the dispersion seen in the middle of the period distribution. Similarly, Fig. 10 shows that the $\dot{P}$ distributions perform worse near the upper tail. The figure also highlights the deviation seen in the CDFs for $\dot{P}$ in the lower end. The $l$ QQ plot is the least interesting since it shows that all the models are in reasonable agreement with the observations since they all lie close to the diagonal line. On the other hand, the $b$ QQ plot has an S-shaped distribution which clearly indicates that the model distributions have heavy tails. If any of these models is assumed to be an acceptable representation of the real sample, than these heavy tails can be interpreted as showing the bias towards lower latitudes in the observed sample. This would mean that there are more MSPs waiting to be discovered further away from the Galactic plane.
@@ 5.2 Canonical pulsars (CPs)
Figure 12: CDF of observable quantities for synthesized MSPs with the 2D-FP $\gamma-$ray Luminosity model compared with the 141 observed $\gamma-$ray MSPs. The y-axis is normalized to allow better comparison of the distributions between the observed and synthesized sample. In the observed sample, 14 are missing a radio flux and 7 are missing $\gamma-$ray flux and $\dot{E}$ values so they are not included in their respective CDFs. These distributions highlight the parameter space yet to be explored by future searches.
Figure 13: Flux vs DM distribution for synthesized MSPs with 2D-FP $\gamma-$ray luminosity model and the heuristic luminosity model for CPs. The red dotted line corresponds to the 100 $\mu$Jy sensitivity limit achieved by targeted radio searches at L-band by TRAPUM using MeerKAT.
@@ 5.3 Predictions for future radio surveys of PLUIDs
The synthesized samples for both MSPs and CPs can be used to make predictions for future targeted radio searches of PLUIDs by looking at sources which have radio fluxes above the sensitivity limit of the deepest searches done so far. We use the sensitivity limit from the searches done by Clark et al. (2023) using MeerKAT, which is 100 $\mu$Jy. In Fig. 13, for MSPs we see that most of the Fermi-LAT detectable MSPs fall below this sensitivity, as expected from the flux distribution in Fig. 12. There are only 143 sources which are above this limit, slightly more than the 138 known radio-loud $\gamma-$ray MSPs. This suggests that, according to this model, most of the Fermi-LAT detectable MSPs that are loud enough to be detected by targeted searches from current radio telescopes have been discovered. Discovering more MSPs among the PLUIDs would require probing lower radio fluxes. This can be achieved by longer observations using existing instruments or using more sensitive instruments. We show the sensitivity limit of the Five hundred meter Aperture Spherical Telescope (FAST) on this plot as well, which is 10 $\mu$Jy. There are 283 MSPs above this limit. For CPs, we pick the heuristic model as our fiducial model when looking at the DM-Flux distribution in Fig. 13. We see that most of the CPs (137/157) lie above the sensitivity limit. The number of known radio-loud CPs is just 84. This suggests, if the heuristic model is assumed to be a good approximation of $\gamma-$ray luminosity, there are many more radio CPs waiting to be discovered among PLUIDs using existing instruments. This strongly supports future targeted observations of LGL region PLUIDs.
@@
For the MSPs, to further preview the properties of these future discoveries, we show CDFs for the distribution of DM, $\dot{E}$, radio flux at 1400 MHz, and $\gamma-$ray flux for 2D-FP model overlaid with the currently observed MSPs in Fig. 12. For the DM, we see that the model has a longer tail with DMs up to $\sim$400 pc cm-3 whereas the maximum DM in the observed sample is 161 pc cm-3. This is indicative of the bias towards MSPs closer to Earth, given DM is a proxy for distance. As can be seen in the sky distribution in Fig. 14, there are still a lot of MSPs to be discovered among the PLUIDs in the HGL region. This is especially important for targeted radio searches since pulsars away from the Galactic plane are not only easier to detect due to mitigated levels of scintillation and scattering but also easier to observe since time on radio telescopes is generally easier to obtain due to such regions being away from the highly oversubscribed sidereal time ranges of most observatories. The remaining three distributions shown in Fig. 12 for the spin-down energy ($\dot{E}$), radio flux and $\gamma-$ray flux peak at lower values than the observed sample showing that the observed sample is biased towards high $\dot{E}$ sources and/or those closer to Earth. For CPs, as shown in Fig. 14, future discoveries are expected primarily in the LGL region.
@@ 6 Conclusions
For CPs, all the models tested satisfy the number constraints. However, the two models adopted from the literature produce an inadequate resemblance to the observed sample when distributed in the $P-\dot{P}$ plane. This trend continues when looking at the distribution of other observables as shown in Fig. 11. In order to get any sort of agreement in $P-\dot{P}$ space, we had to invoke an abitrary cut-off in the $\gamma-$ray beaming fraction to eliminate model pulsars with $\dot{E}<10^{32}$erg s-1. This is a surprising result, especially since, as can be seen in Fig. 3, the radio synthesis generally does an excellent job at reproducing the observed radio population. Therefore, more work is need to understand the $\gamma-$ray emission in CPs. Future work focused on improving the FP models would also benefit from accounting for larger than the observed sample ages seen in the synthesized sample. Our work also shows that the $\gamma-$ray beaming model of Song et al. (2024) is incompatible with the FP relations since the majority of CPs produced by them have $\dot{E}$s too low to be affected by beaming.
@@
We thank David Smith and Simon Johnston for insightful discussions and for sharing code to carry out the $\gamma-$ray flux selection (Smith) and the radio pulsar $P-\dot{P}$ modeling (Johnston). We acknowledge support from a graduate stipend provided by the Eberly College of Arts and Sciences at West Virginia University.

### v2

@@ 4 Results
We evaluate different models based on two main approaches. First, we simply look at the number of Fermi-LAT detectable $\gamma-$ray pulsars predicted by each model and compare it with the known number of pulsars plus the number of PLUIDs in the 4FGL-DR4 Fermi-LAT catalog. Second, we look at the distribution of the modeled sample’s period ($P$), period derivative ($\dot{P}$), Galactic longitude, $(l)$ and latitude $(b)$ distributions to see how they compare with the observed sample. We do this as a three step qualitative process by first looking at the distribution in the $P-\dot{P}$ plot, then comparing the parameters’ cumulative distribution functions (CDFs) and finally evaluating their quantile-quantile (QQ) plots (Wilk & Gnanadesikan, 1968). For a given CDF, every probability value $p$ from the ordinate can be mapped to a quantile value $q(p)$ on the abscissa. For a particular parameter, plotting the $q(p)$ function for two different distributions (in this case, observed CDFs versus model CDFs of the parameter) on the same two-dimensional plot results in a QQ plot for that parameter. If both the distributions are identical, the resulting plot would be a straight line with a gradient of one, passing through the origin. More importantly, if either of the two distributions is a linear function of the other, the resulting QQ plot would still be a straight line albeit with a different slope and $y$-intercept. Any differences between the two distributions, specially in their tails, can be easily spotted on a QQ plot as deviations from linearity. These deviations are easier to spot then noticing subtle changes in the CDFs themselves which is why QQ plots were employed as a visual statistical tool in this work. The purpose of these investigations was not to select an overall preferred model, but rather to check how the underlying distribution resulting from each model compares with the observed one for a given parameter. We also show the $P-\dot{P}$ plot for the resulting populations from our radio synthesis compared with the observed sample which was used to run the Monte Carlo simulation in Fig. 3.
@@ 4.1 Millisecond pulsars (MSPs)
We show the number of Fermi-LAT detectable $\gamma-$ray pulsars predicted by each luminosity model in Fig. 2, along with the acceptable region described in Section 2 highlighted in green. We compare the observed and synthesized distribution for four of these parameters by looking at CDFs shown in Figs. 5, 6, 7 and 8. The observed sample contains 141 $\gamma-$ray MSPs which are in the Galactic plane (i.e., excluding globular cluster MSPs J1823--3021A and J1824--2452A) except for the $\dot{P}$ CDFs in 6. The observed sample in $\dot{P}$ CDFs contains 134 MSPs because 7 MSPs do not have a recorded $\dot{P}$ values in 3PC. Finally, we also provide a $P-\dot{P}$ plot for each of the models in Fig. 4. In this figure, each $P-\dot{P}$ plot has the same number of pulsars. This was done by randomly sampling 134 pulsars to match the number of known pulsars which have measured $\dot{P}$ and are shown in the same figure.
@@
Figure 5: CDFs showing the distribution of period for the observed and model populations of MSPs. The red, solid line in each plot represents the observed population and is labeled as “Obs” in the legend. The dashed, black line represents the model population and is labeled as “Mod” in the legend.
Figure 6: CDFs showing the distribution of period derivative for the observed and model populations of MSPs. The red, solid line in each plot represents the observed population and is labeled as “Obs” in the legend. The dashed, black line represents the model population and is labeled as “Mod” in the legend.
Figure 7: CDFs showing the distribution of Galactic longitude for the observed and model populations of MSPs. The red, solid line in each plot represents the observed population and is labeled as “Obs” in the legend. The dashed, black line represents the model population and is labeled as “Mod” in the legend.
Figure 8: CDFs showing the distribution of Galactic latitude for the observed and model populations of MSPs. The red, solid line in each plot represents the observed population and is labeled as “Obs” in the legend. The dashed, black line represents the model population and is labeled as “Mod” in the legend.
@@ 4.2 Canonical pulsars (CPs)
The 3D-plane suggests 440 Fermi-LAT detectable $\gamma-$ray sources, the 2D-plane suggest 257 of them and the heuristic model, which included beaming, suggests 157 of them. We show the $P-\dot{P}$ distribution for CPs compared with the observed sample in Fig. 9. Since the number of predicted CPs is more than the observed sample for all three models, we chose to show only 150 randomly sampled CPs for each model in these plots. We also show the number of radio-loud $\gamma-$ray sources in this plot. We define radio-loud as the CP having a radio flux of greater than 100 $\mu$Jy.
@@ 5 Discussion
It is important to assess that the radio synthesis results in a population compatible with the observed radio pulsars. This can be done by qualitatively using the $P-\dot{P}$ plot shown in Fig. 3. We can see that the CP synthesis does well, as expected from the work of Johnston & Karastergiou (2017). Similarly, the MSP synthesized sample is able to cluster around the same point as the observed sample. The main discrepancy between the model and observed pulsars in Fig. 3 is the lack of synthesized pulsars in the 20–100 ms period range. This is not surprising given that our modeling does not account for high-mass binary evolution which is thought to produce most of the pulsars in this region (i.e., the double neutron star binaries and the disrupted recycled pulsars; Tauris & van den Heuvel, 2023). These systems are not expected to have a significant impact on our main findings.
@@ 5.1 Millisecond pulsars (MSPs)
[…]
We can find further insights into these distributions by looking at how each quantile of the parameter distributions compares with each other in a QQ plot. For MSPs, this is shown in Fig. 10. All $\gamma-$ray luminosity models except for OG produce high period ($P>9$ ms) MSPs which do not exist in the observed population. Between the $3-5$ ms range, we see a discrepancy between the models. Here, the BPL, LN and 3D FP are more sparse than the observed sample and other models. The $\dot{P}$ distributions in Fig. 10 show that BPL, LN and 3D FP models result in a distribution with longer tails than the observed distribution. All models except BPL and LN do well for rest of the range of $\dot{P}$s until the upper tail. At the upper end, only OG is able to match the observed distribution. The $l$ QQ plot is the least interesting since it shows that all the models are in reasonable agreement with the observations since they all lie close to the diagonal line. On the other hand, the $b$ QQ plot has an S-shaped distribution which clearly indicates that the model distributions have heavy tails. If any of these models is assumed to be an acceptable representation of the real sample, then these heavy tails can be interpreted as showing the bias towards lower latitudes in the observed sample. This would mean that there are more MSPs waiting to be discovered further away from the Galactic plane.
@@ 5.2 Canonical pulsars (CPs)
Figure 12: Flux vs DM distribution for synthesized MSPs with 2D-FP $\gamma-$ray luminosity model and the heuristic luminosity model for CPs. The red dotted line corresponds to the 100 $\mu$Jy sensitivity limit achieved by targeted radio searches at L-band by TRAPUM using MeerKAT. The blue dotted line at 10 $\mu$Jy refers to the approximate sensitivity limit of FAST.
Figure 13: CDF of observable quantities for synthesized MSPs with the 2D-FP $\gamma-$ray Luminosity model compared with the 141 observed $\gamma-$ray MSPs. The y-axis is normalized to allow better comparison of the distributions between the observed and synthesized sample. In the observed sample, 14 are missing a radio flux and 7 are missing $\gamma-$ray flux and $\dot{E}$ values so they are not included in their respective CDFs.
@@ 5.3 Predictions for future radio surveys of PLUIDs
The synthesized samples for both MSPs and CPs can be used to make predictions for future targeted radio searches of PLUIDs by looking at sources which have radio fluxes above the sensitivity limit of the deepest searches done so far. We use the sensitivity limit from the searches done by Clark et al. (2023) using MeerKAT, which is 100 $\mu$Jy. We also use the sensitivity limit of the Five hundred meter Aperture Spherical Telescope (FAST), which is $\sim$10 $\mu$Jy at L-band for a 5 min observation 2025RAA....25a4001H. To make these predictions, we select a fiducial model as well. For MSPs, the main constraint that we place for this selection is that the model should be within the green region of Fig. 2. From the three models that satisfy this requirement, we make a decision based on the qualitative analysis of the $P-\dot{P}$ distribution to use 2D FP as our fiducial MSP model. In the MSP panel of Fig. 12, we see that most of the Fermi-LAT detectable MSPs fall below the MeerKAT sensitivity limit There are only 143 sources which are above the MeerKAT limit, slightly more than the 138 known radio-loud $\gamma-$ray MSPs. This suggests that, according to the 2D FP model, most of the Fermi-LAT detectable MSPs that are loud enough to be detected by the most sensitive targeted searches that have been done so far have been discovered. There are 283 MSPs above the FAST sensitivity limit. This suggests that targeted searches of PLUIDs by FAST could potentially result in discovery of $283-143=140$ MSPs, doubling the current sample size. Discovering more MSPs among the PLUIDs would require probing even lower radio fluxes. This can be achieved by longer observations using existing instruments or by using more sensitive instruments such as the upcoming Square Kilometer Array telescope. For CPs, we pick the heuristic model as our fiducial model due to the discussion about the FP models in section 5.2. Looking at the DM-flux distribution in Fig. 12, we see that most of the CPs (137/157) lie above the sensitivity limit of MeerKAT. All of them are above the limit of FAST. The number of known radio-loud CPs is just 84. This suggests, if the heuristic model is assumed to be a good approximation of $\gamma-$ray luminosity, there are many more radio CPs waiting to be discovered among PLUIDs using existing instruments. This strongly supports future targeted observations of LGL region PLUIDs.
@@
For the MSPs, to further preview the properties of these future discoveries, we show CDFs for the distribution of DM, $\dot{E}$, radio flux at 1400 MHz, and $\gamma-$ray flux for the 2D-FP model overlaid with the currently observed MSPs in Fig. 13. For DM, we see that the model has a longer tail with DMs up to $\sim$400 pc cm-3 whereas the maximum DM in the observed sample is 161 pc cm-3. This is indicative of the bias towards MSPs closer to Earth, given DM is a proxy for distance. As can be seen in the sky distribution in Fig. 14, the 2D FP model suggests that there are still a lot of MSPs to be discovered among the PLUIDs in the HGL region. This is especially important for targeted radio searches since pulsars away from the Galactic plane are not only easier to detect due to mitigated levels of scintillation and scattering but also easier to observe since time on radio telescopes is generally easier to obtain due to such regions being away from the highly oversubscribed sidereal time ranges of most observatories. The remaining three distributions shown in Fig. 13 for the spin-down energy ($\dot{E}$), radio flux and $\gamma-$ray flux peak at lower values than the observed sample showing that the observed sample is biased towards high $\dot{E}$ sources and/or those closer to Earth.
@@ 6 Conclusions
For CPs, all the models tested satisfy the number constraints. However, the two models adopted from the literature produce an inadequate resemblance to the observed sample when distributed in the $P-\dot{P}$ plane. This trend continues when looking at the distribution of other observables as shown in Fig. 11. In order to get any sort of agreement in $P-\dot{P}$ space, we had to invoke an abitrary cut-off in the $\gamma-$ray beaming fraction to eliminate model pulsars with $\dot{E}<10^{32}$erg s-1. This is a surprising result, especially since, as can be seen in Fig. 3, the radio synthesis generally does an excellent job at reproducing the observed radio population. Therefore, more work is needed to understand the $\gamma-$ray emission in CPs. Future work focused on improving the FP models would also benefit from accounting for larger than the observed sample ages seen in the synthesized sample of CPs. Our work also shows that the $\gamma-$ray beaming model of Song et al. (2024) is incompatible with the FP relations since the majority of CPs produced by them have $\dot{E}$s too low to be affected by beaming.
@@
We thank David Smith and Simon Johnston for insightful discussions and for sharing code to carry out the $\gamma-$ray flux selection (Smith) and the radio pulsar $P-\dot{P}$ modeling (Johnston). We also thank the anonymous referee for their valuable feedback which helped improve this manuscript. We acknowledge support from a graduate stipend provided by the Eberly College of Arts and Sciences at West Virginia University.